Probability Formulae (AQA A-Level Mathematics): Flashcards

📚Flashcards
Probability Formulae
Sign up to keep revising.Create a free account to study more flashcards and track your progress.

Practise the cards

10 cards from this deck

Show

Basic probability formula for event AA

P(A)=FavourableTotalP(A) = \frac{\text{Favourable}}{\text{Total}}

Complementary probability formula

P(A′)=1−P(A)P(A') = 1 - P(A)

Addition rule for P(A∪B)P(A \cup B)

P(A∪B)=P(A)+P(B)−P(A∩B)P(A \cup B) = P(A) + P(B) - P(A \cap B)

Multiplication rule for independent events AA and BB

P(A∩B)=P(A)×P(B)P(A \cap B) = P(A) \times P(B)

Multiplication rule for dependent events AA and BB

P(A∩B)=P(A)×P(B∣A)P(A \cap B) = P(A) \times P(B \mid A)

Conditional probability P(A∣B)P(A \mid B) formula

P(A∣B)=P(A∩B)P(B)P(A \mid B) = \frac{P(A \cap B)}{P(B)}

Meaning of P(B∣A)P(B \mid A)

Probability of BB given AA has occurred

Total probability formula using conditional prob

P(A)=∑P(Bi)×P(A∣Bi)P(A) = \sum P(B_i) \times P(A \mid B_i)

Bayes' theorem formula

P(Bi∣A)=P(Bi)×P(A∣Bi)P(A)P(B_i \mid A) = \frac{P(B_i) \times P(A \mid B_i)}{P(A)}

Probability of at least one event occurring

P(At least one)=1−P(None)P(\text{At least one}) = 1 - P(\text{None})

Explore AQA A-Level Mathematics Revision Notes by Topics

Explore AQA A-Level Mathematics Model Answers by Topics

Explore AQA A-Level Mathematics Quizzes by Topics

Explore AQA A-Level Mathematics Exam Questions by Topics

Join 100,000+ A-Level students studying Flashcards with us.

Select your subjects, and get access to A+ resources today.